# Leveraging Large Language Models for Explainable Activity Recognition in Smart Homes: A Critical Evaluation

### v1

@@ 4.3.1. Pre-processing and dataset splitting for ADL recognition
4.3.2. User Surveys
To assess the quality of the explanations from the point of view of non-expert users, we performed two user-based surveys: one for the MARBLE dataset and another for the Home B of the UCI ADL dataset. We excluded Home A from this analysis since it is a significantly simpler dataset compared to Home B (as also reflected by our results in Section 4.4). Frequently, in this dataset there is a only predominant state in the time windows, leading to relatively simplistic explanations.
@@ 4.4.1. Recognition rate
Tables 3 and 4 compare the recognition rate of ADL-LLM and LLMe2e measured by the weighted F1-score [^3] For the sake of fairness, we did not consider dynamic manipulations as input for LLMe2e since these were not considered in the ADL-LLM paper (Civitarese et al., 2024).. The results show that LLMe2e achieves a similar recognition rate for both datasets despite the lower-level input representation (JSON format) and the additional task of generating explanations.
@@ 5.1. Over-reliance
Considering the methods proposed in this paper, LLMe2e has a complete knowledge of the input data and of the ADL classification process to generate explanations, potentially offering a more accurate explanation of the decision process. On the other hand, LLMExplainer only has to generate an explanation given the most important features obtained from a data-driven classifier. We observed that LLMExplainer, in some instances, tends to generate explanations that are linguistically good but not supported by the observed sensor data. Hence, even though LLMExplainer produces appealing explanations highly rated in our surveys, these can be sometimes misleading, fostering an illusion of certainty and reliability. This problem becomes relevant when considering wrong predictions, as illustrated in the examples of Table 5.
@@ 5.3. Applicability
5.4. Cost and privacy issues
@@
In our experimental evaluation, we used a third party cloud-based LLM that is known to perform very well on many tasks beyond ADL recognition. However, the adoption of these models in real deployments may have two drawbacks: costs and privacy issues.
Considering the costs, both LLMe2e and LLMExplainer require a remote call to the LLM APIs for each window. At the time of writing, a single window for LLMe2e is costing on average $0.0085\$$ when using gpt-4o. Considering windows of $16$ seconds with $0.8$ of overlap, as we did for the MARBLE dataset, operating the system continuously would cost approximately $230\$$ per day.
Open-source LLMs are a promising alternative to reduce these costs. However, the most accurate models still require powerful computing machines equipped with GPUs and substantial amounts of RAM, resulting in costs not only for the hardware, but also for energy consumption. Dedicated machines could be deployed in a telemedicine infrastructure to optimize the costs by performing ADL recognition for a wide number of subjects. It is unclear if, in the future, more lightweight language models (e.g., small models like Phi-4 (Abdin et al., 2024) or quantized versions of open-source models) will potentially run on smart home gateways with acceptable performance and accuracy. Open-source models could also mitigate privacy problems since data about subjects’ actions in the home and their activities could be processed by a trusted entity and not released to an untrusted third-party provider.
@@ 6. Conclusion and Future Work
Among the many research directions that this work is opening, we plan to test open-source LLM-based solutions on a large unlabeled dataset that we are acquiring from smart homes as part of a collaboration project with neurologists [^4] https://ecare.unimi.it/pilots/serenade/.

### v2

@@ 4.3.1. Pre-processing and dataset splitting for ADL recognition
4.3.2. Quantitative Evaluation
The experimental evaluation of our approaches consists of quantitatively measuring both recognition rate and explanation quality. Considering the recognition rate, we use standard measures like the weighted F1 scores and confusion matrices.
On the other hand, a quantitative assessment of the explanations’ quality from the point of view of non-expert subjects requires user-based studies usually adopted in Human Computer Interaction (HCI) works. Indeed, XAI is human-centered, as its primary goal is to make AI decisions understandable and trustworthy to humans. Consequently, evaluating explanations quality inherently requires human judgment  (Mohseni et al., 2021).
Hence, we performed two user-based surveys: one for the MARBLE dataset and another for the Home B of the UCI ADL dataset. We excluded Home A from this analysis since it is a significantly simpler dataset compared to Home B (as also reflected by our results in Section 4.4). Frequently, in this dataset there is a only predominant state in the time windows, leading to relatively simplistic explanations.
@@ 4.4.1. Recognition rate
Tables 3 and 4 compare the recognition rate of ADL-LLM and LLMe2e measured by the weighted F1-score [^3] For the sake of fairness, we did not consider dynamic manipulations as input for LLMe2e since these were not considered in the ADL-LLM paper (Civitarese et al., 2025).. The results show that LLMe2e achieves a similar recognition rate for both datasets despite the lower-level input representation (JSON format) and the additional task of generating explanations.
@@ 5.1. Over-reliance
Considering the methods proposed in this paper, LLMe2e has complete knowledge of the input data and of the ADL classification process to generate explanations, potentially offering a more accurate explanation of the decision process. On the other hand, LLMExplainer only has to generate an explanation given the most important features obtained from a data-driven classifier. We observed that LLMExplainer, in some instances, tends to generate explanations that are linguistically good but not supported by the observed sensor data. Hence, even though LLMExplainer produces appealing explanations highly rated in our surveys, these can be sometimes misleading, fostering an illusion of certainty and reliability. This problem becomes relevant when considering wrong predictions, as illustrated in the examples of Table 5.
@@ 5.3. Limitations in setups dominated by PIR sensors
5.4. Privacy issues, financial cost, and scalability
@@
In our experimental evaluation, we used a third-party cloud-based LLM that is known to perform very well on many tasks beyond ADL recognition. However, the adoption of these models in real deployments may have three drawbacks: privacy issues, financial cost, and scalability.
Considering privacy, personal data about human activities is sensitive, as it can reveal sensitive personal habits and medical conditions. Hence, all data acquired in the smart home should be transmitted, processed, and stored with state-of-the-art security methods (Bettini et al., 2025). In particular, outsourcing such information to untrusted third-party providers (e.g., LLM services) exposes users to potential privacy issues. Deploying an open-weights LLM on a private machine of a trusted entity in a trusted domain is a promising alternative to mitigate privacy problems. For instance, considering medical applications, dedicated servers could be deployed in a telemedicine infrastructure. It is also possible that, in the future, more lightweight language models (e.g., small models like Phi-4 (Abdin et al., 2024) or quantized versions of open-source models) will potentially run on smart home gateways with an acceptable trade-off between performance and accuracy.
However, to reduce privacy risks using open-weights models, a careful configuration of security settings is required. This requires, for example, encrypting data on-transit and at-rest. From the security point of view, poorly configured servers may also increase the risks of unauthorized access and/or model replacements with tampered or backdoored versions that can behave maliciously. Moreover, LLMs can be tricked into executing malicious commands or generating harmful outputs if not properly sandboxed.
Considering financial cost, both LLMe2e and LLMExplainer require a remote call to the LLM APIs for each window. At the time we performed our experiments using the cloud-based gpt-4o model, a single window for LLMe2e was costing on average 0.0085\$0.0085\\$. Considering windows of $16$ seconds with $0.8$ of overlap, as we did for the MARBLE dataset, operating the system continuously at this price would cost approximately 230\$230\\$ per day, which is a very high cost. Despite the price for LLM usage is likely to decrease significantly with the design of more efficient and powerful models, this is an issue to consider. On the other hand, even local deployments of open-weights models incur financial costs. Indeed, the most accurate open-weights LLMs currently still require powerful servers, resulting in costs not only for the hardware, but also for energy consumption. In this scenario, a telemedicine platform could optimize the costs by performing ADL recognition for a large number of subjects.
Finally, considering scalability, using third-party models introduces some bottlenecks to take into account, especially considering future smart homes that will likely be equipped with a significantly high number of sensing devices generating large inputs for the LLM. Indeed, third-party services usually have usage limits, for example in the number of requests, the number of tokens processed per minute, and the number of tokens for each request. To maximize throughput, implementing parallel API calls also requires adopting strategies like retries with exponential backoff, concurrency control, and request queuing. Another factor impacting scalability is the latency per request, that is affected by network communication delays. On the other hand, to achieve the necessary scalability with open-weights models, the telemedicine platform should run multiple instances of the LLM across multiple machines with GPUs. This setup, with the adequate hardware resources, is crucial for potentially handling the many concurrent requests coming from numerous subjects, while also ensuring high availability and fault tolerance in case of individual instance failures. Considering future smart homes equipped with numerous sensing devices, we believe that recent open-weight models already allow processing inputs that are sufficiently large (e.g., Gemma 3 27b has a context window of $128k$ tokens). As an alternative, distributing lightweight language models across the homes directly in edge devices (e.g., in the gateway) would further improve scalability. Hence, while open-weight models represent a promising direction, their deployment should be properly designed to take costs, scalability, and security configurations into account.
5.5. Generalization and Personalization of LLMe2e
Our LLMe2e approach leverages LLMs pretrained on vast and diverse textual corpora to perform zero-shot activity recognition without requiring labeled data. As demonstrated in our experiments across multiple datasets, LLMe2e achieves high recognition rates in home environments that differ in layout, object configurations, and user behavior.
However, the generalization capabilities of LLMe2e may still be limited considering individualized behavioral patterns, culturally specific routines, or unique home configurations not captured by the general LLM knowledge.
To mitigate this challenge for the smart home environments, LLMe2e system prompt requires details about the home infrastructure and the set of monitored objects (see Figure 3). These cues guide the LLM in interpreting sensor data for ADLs recognition. As future work, we plan to expand the system prompt to incorporate user-specific context—such as habitual routines, social or cultural norms, and personal schedules to further personalize ADLs recognition.
While prompt engineering offers a lightweight method for adaptation, it may not suffice in all cases. Fine-tuning the LLM on labeled data from the target environment could, in principle, yield better personalization with individual users. However, such an approach is often constrained by the high costs, privacy risks, and logistical complexity of acquiring labeled data in real-world smart home deployments.
To address these limitations, semi-supervised learning strategies present a promising avenue. Techniques that combine unsupervised clustering with active learning (e.g., (Hiremath et al., 2022; Arrotta et al., 2023)) could enable minimal supervision by selecting only the most informative examples for labeling from the massive amount of unlabeled sensor data.
However, it is important to note that LLMe2e goes beyond classification; it also generates natural language explanations, which are the core focus of this paper. These explanations do not have ground-truth supervision and depend on the LLM’s generative capabilities. Fine-tuning only the ADLs recognition task may risk degrading explanation quality. To the best of our knowledge, no fine-tuning method currently guarantees preservation of generative reasoning alongside predictive accuracy.
As an alternative, few-shot prompting offers a promising strategy for personalization. As also demonstrated in ADL-LLM (Civitarese et al., 2025), by including a small number of labeled examples directly into the prompt, the model can be guided to improve both prediction and explanation without altering its parameters. This approach maintains the integrity of the generative capabilities while enabling user-specific adaptation.
Hence, we will explore hybrid personalization strategies integrating few-shot prompting with semi-supervised learning techniques, enabling LLMe2e to adapt more effectively to individual users.
This future line of research also opens up opportunities for long-term adaptation. Indeed, by continuously observing unlabeled data, it becomes possible to detect shifts in an individual’s activity patterns over time. Integrating concept drift detection with active learning would allow the system to selectively update the labeled examples used in few-shot prompting, ensuring that the model remains aligned with the subject’s evolving routines and lifestyle.
@@ 6. Conclusion and Future Work
As part of a collaboration project with neurologists  [^4] https://ecare.unimi.it/pilots/serenade/, we are collecting a large unlabeled dataset from smart homes of individuals with Mild Cognitive Impairment (MCI). Hence, in future work, we plan to test our methods on this dataset, possibly also exploring open weights Large and Small Language Models.
Acknowledgements.
This work was supported in part by projects MUSA and FAIR under the NRRP MUR program funded by the EU-NGEU. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the Italian MUR. Neither the European Union nor the Italian MUR can be held responsible for them.